OpenStudy (anonymous):

A block is placed on an inclined plane. The static coefficient of friction between the block and the plane is .300. What is the maximum angle the plane can have without the block starting to slide?

OpenStudy (shamim):

Ket the angle between the inclined plane n horizontal plane is theta

OpenStudy (shamim):

*let

OpenStudy (shamim):

Let the mass of the block is m

OpenStudy (shamim):

So weight of mass m =mg

OpenStudy (shamim):

Component of weight mg along inclined plane=mg*sin theta

OpenStudy (anonymous):

friction(.300) + mgsin(theta) =ma?

OpenStudy (shamim):

Component of weight mg perpendicular to the inclined plane=mg*cos theta

OpenStudy (anonymous):

oh so .300 + mgcos(theta)=0

OpenStudy (shamim):

Acceleration a=0 in ur equation. Because the block is static

OpenStudy (anonymous):

yes i forgot that

OpenStudy (shamim):

So Frictional force=mg*sin theta Right?

OpenStudy (anonymous):

so .300=mgcos(theta)

OpenStudy (anonymous):

yes

OpenStudy (shamim):

No

OpenStudy (anonymous):

mgcos(theta)

OpenStudy (shamim):

Frictional force=coefficeint of static friction*mg*cos theta

OpenStudy (shamim):

R=mg*cos theta

OpenStudy (anonymous):

what is r?

OpenStudy (shamim):

Frictional force=0.3*mg*cos theta

OpenStudy (shamim):

R is a force acting on the block by the inclined plane

OpenStudy (anonymous):

oh ok

OpenStudy (shamim):

mg*cos theta Is a force acting on the inclined plane by the block

OpenStudy (anonymous):

so how do you get a value if you only have 1 to plug in

OpenStudy (shamim):

Ok 0.3*mg*cos theta=mg*sin theta Theta=?

OpenStudy (shamim):

Anyway i m frm an android cell phone

OpenStudy (shamim):

So its not possible to draw

OpenStudy (shamim):

Can u draw it

OpenStudy (anonymous):

ok ill try

OpenStudy (anonymous):